In the kingdoms of Freedonia and Socialvania the production function is y = Ak^(2/3)h^(1/3). In both kingdoms the fraction of output saved is 15% the depreciation rate is 3% and the population growth rate is zero. Assume A is equal to one.

(a) What is the ratio of steady-state level of output per worker?

( b) What is the degree to which schooling differences explain differences in income per capita? Assume average schooling in Freedonia is 9 years and average schooling in Socialvania is 6 years.
Answer #1

a) Production function y = f (K, L) is output per worker.
In this case y = Ak^(2/3)h^(1/3) = 1*0.15^(2/3)*.03^(1/3) = .189
He h =.03
k= .15
so the ratio of output per worker is .189
b) Education helps in generating quality human capital and improves productivity. This in turn increases output per capita or income per capita. Research shows if there is a rise of 1 percent in GDP per capital for a duration of five years, then it corresponds to a rise of .3% in number of years of schooling on average.